atomic mass of boron is 10.81.it has two isotopes with 80% and 20% abundance respectively. the atomic mass of the isotopes havinf 80% abundance is 11.01.the atomic mass of the other isotope is

Grade:11

1 Answers

Sheetal kumari
6 years ago
Atomic mass of boron is 10.81.percentage abundance of both isotopes are given .Atomic mass of one isotope is 11.01then[%abundance×mass of isotope+%abundance×mass of other isotope]÷100=atomic mass of element[80×11.01+20×x]÷100=10.81880.8+20x=1081x comes to be 10.01hope it helps you